OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[\frac{ 15x + 15 }{ x^{2}- 1}\]

OpenStudy (anonymous):

@whpalmer4 Could you help me? :/

OpenStudy (whpalmer4):

Factor numerator and denominator and cancel common factor(s)

OpenStudy (anonymous):

(15x+15)/x^2-1=15(x+1)/(x+1)(x-1)=15/x-1

OpenStudy (whpalmer4):

Numerator becomes 15(x+1) Denominator is difference of squares so factors to (x-1)(x+1)

OpenStudy (anonymous):

Answers: -15 15/x-1 30/x-1 15

OpenStudy (anonymous):

So B is correct?

OpenStudy (anonymous):

yes,itz b

OpenStudy (anonymous):

Thank youu!